OpenStudy (anonymous):

Solve the logarithmic equation for x. 1.) log(x − 8) = 3 x= ? 2.) 3 − log(6 − x) = 2 x=?

OpenStudy (callisto):

@sorour \[\log (\frac{x}{y}) = \log (x) - \log (y)\] Also, do not give a full solution.

OpenStudy (anonymous):

Uh so what's the final answer?

OpenStudy (anonymous):

yeah you are right,sorry

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I just need an answer lol

OpenStudy (anonymous):

I've done all my work already. I'm trying to do a practice test for my test that is today and I need answers to check if I'm right. That's all

OpenStudy (anonymous):

I don't feel like going through the whole process over again sorry

OpenStudy (callisto):

It would be better if you can show us your work so that we can check it for you.